According to the following equation, how many moles of aluminum are needed to form 3.7 mol Al2O3?

4Al + O2
1
Expert's answer
2020-11-11T10:35:01-0500

Solution:

The balanced chemical equation:

4Al + 3O2 = 2Al2O3

According to the chemical equation:

n(Al)/4 = n(Al2O3)/2

n(Al)/2 = n(Al2O3)

n(Al) = 2 × n(Al2O3)

Hence,

n(Al) = 2 × 3.7 mol = 7.4 mol

Moles of Al = 7.4 moles


Answer: 7.4 moles of aluminum (Al) are needed to form 3.7 mol Al2O3